# Malsight: Exploring Malicious Source Code and Benign Pseudocode for Iterative Binary Malware Summarization

### v1

@@ Abstract
To this end, we propose Malsight, a novel code summarization framework that can iteratively generate descriptions of binary malware by exploring malicious source code and benign pseudocode. Specifically, we construct the first malware summaries, MalS and MalP, using an LLM and manually refine this dataset with human effort. At the training stage, we tune our proposed MalT5, a novel LLM-based code model, on the MalS dataset and a benign pseudocode dataset. Then, at the test stage, we iteratively feed the pseudocode functions into MalT5 to obtain the summary. Such a procedure facilitates the understanding of pseudocode structure and captures the intricate interactions between functions, thereby benefiting the usability, accuracy, and completeness of summaries. Additionally, we propose a novel evaluation benchmark, BLEURT-sum, to measure the quality of summaries. Experiments on three datasets show the effectiveness of the proposed Malsight. Notably, our proposed MalT5, with only 0.77B parameters, delivers comparable performance to much larger ChatGPT3.5.
@@
1 Introduction
The AV-TEST Institute [1] recently reported that over 450,000 new malicious files and potentially unwanted applications are registered daily, showing a high demand for malware understanding. Binary malware summarization [2] is a reverse engineering [3] task that aims to automatically generate concise human-readable descriptions of binary executable malicious files. The summarization provides security analysts with a quick understanding of the malware’s functionality and patterns when source code is unavailable, thereby benefiting a wide range of applications such as malware cracking [4] [5] [6], malware family classification [7] [8], binary code similarity detection [9] [10] [11], and large-scale malware behavior analysis [12] [13] [14].
@@ 1 Introduction
Figure 1: The comparison of source code (left) and its pseudocode (right). The pseudocode includes significantly more content and a more complex structure, and it also strips key semantic cues such as function names.
@@
Existing reverse engineering tools, such as IDA [15] and Ghidra  [16], can decompile executables into higher-level C-like pseudocode, while they still lack easy-to-understand semantics information. Consequently, a line of efforts attempts to generate human-readable summaries based on pseudocode. Early studies rely on manual parsing or rule-based summary generation [17] [18]. Recent large language models (LLMs), such as BinT5 [19], HexT5 [20], CodeGen [21], and WizardCoder [22], have shown great potential to produce more informative summaries. However, these data-driven approaches still face critical issues, including poor usability, inaccurate explanations, and inaccurate explanations [2]. Figure 1 shows the underlying reasons for the above issues by comparing the source code of the function “initLevel” to the corresponding pseudocode. We observe that the pseudocode presents 1) significantly more content that increases from 20 lines in source to 117 lines in the pseudocode, 2) a more complex and obscure structure with multi-level nesting and entangled logic. The pseudocode involves 29 more calls and 29 more if statements compared to the source code at the left, 3) stripping key semantic cues such as variable names and function names. For example, the function “initLevel” in source code is transferred to a meaningless symbol “sub_404018”.
@@
To address the above challenges, we present Malsight, a novel binary malware summarization framework that can iteratively generate descriptions of executable malware by exploring malicious source code and benign pseudocode. The proposed Malsight involves three key ingredients, including a malware dataset MalS, an LLM-based malware summarization model MalT5, and an evaluation metric BLEURT-sum. We describe the workflow of the proposed framework in four steps as follows.
@@
Constructing MalS: As an LLM-based summarization model heavily relies on high-quality annotations to learn to align with domain-specific knowledge, it necessitates high-quality malware pseudocode summaries to fine-tune the LLM. However, the public malware pseudocode summarization dataset is unavailable so far, and building such a benchmark is quite challenging as it requires huge human involvement for accurate annotations. Figure 1 illustrates three challenges of understanding malware pseudocode. To tackle this issue, we alternatively construct MalS, a large-scale summarization dataset using an LLM model, and malicious C language source code crawled from GitHub. The proposed MalS involves nearly 90,000 malware source functions, with 20 types of malware functions. We also construct a small dataset MalP for testing. We detail such a procedure in Section 4.3.
@@
Training MalT5: We use CodeT5+ [23] as the foundation model of our MalT5. We sequentially fine-tune the proposed MalT5 model on the MalS dataset and an existing benign pseudocode summarization dataset [19]. The underlying intuition is that the malicious semantic knowledge from malware source code summarization and function patterns from benign pseudocode summarization, which are learned from the above two datasets, respectively, can be transferred to the generation of malware pseudocode. By doing so, we can properly mitigate the issue of unavailable malware pseudocode summarization datasets. More details are available in Section 4.4.
Performing Generation: We use an existing tool [15] to generate pseudocode of a binary file and then generate summaries using the MalT5 model. We first use IDA to construct the malware call graph and then develop an algorithm to transform the graph into a function list in reverse order. Then we iteratively fed the first function in the list to MALT5 to generate the summary. More details are provided in Section 4.1 and 4.2.
@@
Conducting Evaluation: Previous work [24] indicated that existing metrics for generation tasks, such as Bilingual Evaluation Understudy (BLEU) [25], Metric for Evaluation of Translation with Explicit ORdering (METEOR) [26], Recall-Oriented Understudy for Gisting Evaluation-Longest Common Subsequence (ROUGE-L) [27], may not well-fit for evaluation of the binary malware summarization. We thus employ BLEURT-sum, which is more sensitive to the quality of the pseudocode summary, thereby benefiting the evaluation in practice. More descriptions are given in Section 5.1.
We conduct experiments on three datasets to verify the effectiveness of the proposed Malsight framework for binary malware summarization. The contribution of this paper can be summarized as follows [^1] We will release our Malsight to contribute to the community..
@@
  • A binary malware summarization Framework. We propose Malsight, a novel framework that can iteratively generate descriptions of binary malware by exploring malicious source code and benign pseudocode. Our MalT5 can tackle the challenges of entangled logic and stripped semantics in pseudocode.
@@
  • Large-scale datasets for binary malware summarization. We propose MalS and MalP, two novel datasets that can be used for the LLM training and testing of an LLM of binary malware summarization. To the best of our knowledge, the two datasets are the first in the field, involving nearly 90,000 malicious source functions and 20 types. Our MalS and MalP can serve as a benchmark for various binary malware understanding tasks.
@@
  • An LLM-based binary malware summarization model. We propose MalT5, a novel LLM for the summarization task. The proposed MalT5 is lightweight, with only 0.7B parameters.
@@
  • An evaluation metric for the task: We present BLEURT-sum, a novel evaluation metric that is more sensitive to the quality of pseudocode summarization.
  • Extensive experiments. We conduct extensive experiments on three datasets and provide case studies to show why the proposed framework performs best among all baselines. Results show that our MalT5 achieves comparable performance to ChatGPT3.5.
2 Background
2.1 Malware Analysis Engineering
@@
The field of malware analysis engineering focuses on analyzing the functionality of malware by examining its binaries, typically through static analysis methods that involve observing assembly code or pseudocode [28].
2.1.1 Binary Decompilation
@@
Decompilation [29] converts executable files into human-readable pseudocode [30], which is more concise and structured than disassembled assembly code. Unlike disassembly, which maps instruction encoding directly to assembly statements, decompilation relies on algorithms and patterns (e.g. R2 [31], IDA [15], Ghidra [16]) and emerging methods using LLMs [32]. However, pseudocode lacks semantic information such as function names. Decompiled function names are often unreadable (e.g., sub_4061C0 in IDA Pro) [33], providing a little useful pieces of information for further analysis.
2.1.2 Human Static Analysis
@@
2.2 NLP Technologies
In the Malsight process, we use Bidirectional Encoder Representation from Transformers (BERT) [37] and Text-to-Text Transfer Transformer (T5) [38] architecture language models to complete specific tasks. For the core code summary task, we build a CodeT5+ model combined with transfer learning.
2.2.1 BERT Family
BERT is a large-scale transformer-based language model pre-trained on a wide corpus of text using a self-supervised learning approach. The design of Masked Language Modeling (MLM) and Next Sentence Prediction (NSP) training tasks makes BERT perform well in the tasks of Sequence Labeling (SL), such as Named Entity Recognition (NER).
Building on BERT, CodeBERT [39] learns code semantics through Code-Conditioned masked language modeling (CMLM) and natural language documentation generation (NLG). CodeBERT has been shown to perform well on code-related tasks, and since it is derived from BERT, we have reason to believe that this model can be fine-tuned to solve the problem of SL in pseudocode.
2.2.2 T5 Family
T5 [38], or Text-to-Text Transfer Transformer, is a sequence-to-sequence model based on the Transformer architecture that unifies various Natural Language Processing (NLP) tasks into a single framework, including text classification, question answering, summarization, translation, and text generation.
CodeT5 [40] is an encoder-decoder model supporting code understanding and generation, built on the T5 architecture. It uses Natural Language-Programming Language (NL-PL) bimodal data for pre-training with identifier tagging and masked identifier prediction tasks. CodeT5+ [23] introduces greater architectural flexibility and additional pre-training tasks, with instruction tuning to enhance alignment with natural language instructions. This results in significant performance improvements on various code-related tasks.
Previous works like HexT5 [20] and BinT5 [19] developed datasets to train models for binary code understanding, including code summarization tasks. These efforts demonstrate the potential of T5-based models in binary code summarization for malware analysis.
2.2.3 Transfer Learning
Transfer learning involves training a model on a source task with abundant labeled data to learn general features. When it is difficult to obtain sufficient datasets for training, transfer learning can be used to supplement them with similar or related datasets [41].
In Malsight, we fine-tune the CodeT5+ [23] model to achieve transfer learning from the source code summarization task to the decompiled code summarization task. Besides, we use dynamic and static annotation to implement feature enhancement to compensate for the poor transfer effect caused by the highly limited similarity between the source code and the stripped decompiled code.
2.3 Code Summary Evaluation
In code summary model evaluation, NLP text similarity algorithms compare generated results with a reference test set, replacing costly human evaluations. These algorithms are categorized into word overlap and word embedding measures.
2.3.1 Words’ Overlap Measure
@@
2.3.2 Words’ Embedding Measure
The words’ embedding measure evaluates semantic similarity by analyzing the distance between sentence embeddings in a vector space, often utilizing neural network learning.
@@ 2.3.2 Words’ Embedding Measure
Figure 2: Workflow of Malsight. The procedure involves three steps: graph traversal, annotation generation, and model summary.
3 Motivation and Overview
The construction of the code summary framework mainly includes annotation generation and code summary model construction, as shown in Figure 2. During the evaluation phase, we tested several evaluation methods and found a reasonable way to build an evaluation model for the code summary task. Simultaneously, our work involves the construction of multiple datasets (for subsequent stages of training and evaluation of transfer learning-based models).
3.1 Code Summarization Process
@@
3.1.1 Function List Extraction
As mentioned, existing code summary methods for binary focus only on the internal information of the function. We introduced the call relationship between functions and worked on the entire binary as the processing unit. In other words, when function func_E in Figure 2 calls func_F, it is difficult for the subsequent code summarization model to correctly summarize the functionality of func_E without any information about func_F. (We assume that the function name of func_F has been corrupted.) Constructing a list of reverse call sequential relational functions provides a basis for the subsequent recovery of sub-functions functionality.
3.1.2 Annotation Generation
Iterate through the list of functions (assuming Fun_F has been processed), and Fun_E will first be added with annotations by the static annotator and the dynamic annotator, respectively. Fun_E uses the static annotator to obtain static annotations based on the internal information of the function code, while the dynamic annotator adds dynamic annotations based on the generated summary of the sub-function (Func_F) in the function. The program, in other words, sequentially restores functions according to the Control Flow Graph (CFG) diagram from the outermost to the innermost and passes function summary results inward.
3.1.3 Code LLM Summary
Fun_E(annotated) is then fed into the code summary model for final code summary generation. In our work, we use transfer learning to adapt the model to both the functionality of malware functions and the structural features of decompiled pseudocode. Based on the CodeT5+ model, we have fine-tuned the code summary task. The tokenizer splits the code into tokens and embeddings, incorporating a self-attention mechanism into a complete vector in the encoder. The decoder outputs a fine-tuned prediction based on the code summary.
@@ 3.2 Evaluation Method
Our research has found that existing methods can not simultaneously measure the meaning, structure, word frequency, and other features of the reference sentence and the candidate sentence, so the model’s performance may be misjudged. Taking Figure 3 as an example, two examples show the evaluation results of BLEU, Meteor, and ROUGE-L on two pairs of real code summaries. The figure shows that two code summaries without any semantically related results in high evaluation scores (blue-framed), while two semantically similar code abstracts receive low scores (red-framed), demonstrating the shortage of existing methods. In the following work, we construct an ML-based code summary evaluation method BLEURT-sum by constructing a set of positive and negative samples composed of related sentence pairs and unrelated sentence pairs. We evaluated the usability of the model and prior art by measuring their ability to distinguish between positive and negative samples.
3.3 Datasets Construction
@@
For the two core tasks mentioned above, binary code summary and code summary model evaluation, we build corresponding datasets.
3.3.1 Dataset For Code Summary Model
In order to avoid the data shift problem, the training of the code summary model requires a large dataset of malware pseudocode summary. Unfortunately, malware datasets are typically represented as collections of compiled binary files [48], with the binary code stripped, and possibly structurally confused. Consequently, the creation of code summary datasets for binary malware could be unfeasible without resorting to labor-intensive manual summarization.

### v3

@@ Abstract
To this end, we propose Malsight, a novel code summarization framework that can iteratively generate descriptions of binary malware by exploring malicious source code and benign pseudocode. Specifically, we construct the first malware summary dataset, MalS and MalP, using an LLM and manually refine this dataset with human effort. At the training stage, we tune our proposed MalT5, a novel LLM-based code model, on the MalS and benign pseudocode datasets. Then, at the test stage, we iteratively feed the pseudocode functions into MalT5 to obtain the summary. Such a procedure facilitates the understanding of pseudocode structure and captures the intricate interactions between functions, thereby benefiting summaries’ usability, accuracy, and completeness. Additionally, we propose a novel evaluation benchmark, BLEURT-sum, to measure the quality of summaries. Experiments on three datasets show the effectiveness of the proposed Malsight. Notably, our proposed MalT5, with only 0.77B parameters, delivers comparable performance to much larger Code-Llama.
@@
I Introduction
@@
Figure 1: The comparison of source code (left) and its pseudocode (right). The pseudocode includes significantly more content and a more complex structure and strips key semantic cues such as function names.
The AV-TEST Institute [1] recently reported that over 450,000 new malicious files and potentially unwanted applications are registered daily, showing a high demand for malware understanding. Binary malware summarization [2] is a reverse engineering [3] task that aims to generate concise human-readable descriptions of binary executable malicious files automatically. The summarization provides security analysts with a quick understanding of the malware’s functionality and patterns when source code is unavailable, thereby benefiting a wide range of applications such as malware cracking [4] [5] [6], malware family classification [7] [8], binary code similarity detection [9] [10] [11], and large-scale malware behavior analysis [12] [13] [14].
@@
Existing reverse engineering tools, such as IDA [15] and Ghidra  [16], can decompile executables into higher-level C-like pseudocode while still lacking easy-to-understand semantics information. While tools such as IDA Pro and Ghidra have significantly advanced binary decompilation by transforming machine code into higher-level pseudocode, they still exhibit critical limitations in handling stripped binaries and entangled logic structures. Specifically, these tools often fail to recover semantically meaningful variable names, function identifiers, or accurate control-flow structures when binaries are stripped of symbolic information. This deficiency leads to pseudocode that is not only semantically impoverished but also structurally convoluted, with deeply nested branches and opaque function calls (e.g., sub_404018 instead of initLevel), as demonstrated in Fig. 1. Existing small models, although lightweight, often suffer from insufficient capability to capture the complex semantics and entangled logic in stripped malware pseudocode, leading to incomplete or inaccurate summaries. Consequently, a line of efforts attempts to generate human-readable summaries based on pseudocode. Early studies rely on manual parsing or rule-based summary generation [17] [18]. Recent large language models (LLMs) [19], such as BinT5 [20], HexT5 [21], CodeGen [22], and WizardCoder [23], have shown great potential to produce more informative summaries. However, these data-driven approaches still face critical issues, including poor usability, inaccurate explanations, and inaccurate explanations [2]. Fig. 1 shows the underlying reasons for the above issues by comparing the source code of the function “initLevel” to the corresponding pseudocode. We observe that the pseudocode presents 1) significantly more content that increases from 20 lines in source to 117 lines in the pseudocode, 2) a more complex and obscure structure with multi-level nesting and entangled logic. The pseudocode involves 29 more calls and 29 more if statements compared to the source code at the left, and 3) stripping key semantic cues such as variable names and function names. For example, the function “initLevel” in the source code is transferred to symbol “sub_404018”.
@@
To address the above challenges, we present Malsight  [^1] We will release Malsight’s source code to contribute to the community. , a novel binary malware summarization framework that can iteratively generate descriptions of executable malware by exploring malicious source code and benign pseudocode. The proposed Malsight involves three key ingredients, including a malware dataset MalS, an LLM-based malware summarization model MalT5, and an evaluation metric BLEURT-sum. We describe the workflow of the proposed framework in four steps.
@@
Constructing MalS: LLM-based summarization models rely heavily on high-quality summaries to capture domain-specific knowledge. However, the public malware pseudocode summarization dataset is unavailable so far, and building such a benchmark is quite challenging as it requires substantial human involvement for accurate summarization, due to the challenges outlined in Fig. 1. To address this, we construct MalS, a large-scale dataset with well-constructed summaries based on malicious C-language source code from GitHub. The proposed MalS involves nearly 90,000 malware source functions, with 20 types of malware functions. We also constructed a small dataset, MalP, for testing. We detail such a procedure in Section IV-C.
@@
Training MalT5: We propose MalT5, a model fine-tuned on both the MalS dataset and an existing benign pseudocode summarization dataset [20]. The underlying intuition is that the malicious semantic knowledge from malware source code summarization and function patterns from benign pseudocode summarization, which is learned from the above two datasets, respectively, can be transferred to the summarization of malware pseudocode. On the one hand, this also effectively addresses the challenges brought about by the lengthy content and complex and obscure structure of decompiled pseudo-code mentioned in the previous text. On the other hand, by doing so, we can properly mitigate the issue of unavailable malware pseudocode summarization datasets. More details are available in Section IV-D.
Performing Generation: Initially, IDApro generates pseudocode and extracts the call graph from a binary file. We then develop an algorithm to convert the call graph into a function list, preserving the original call order. This function list is iteratively fed into the MalT5 model to generate summaries, following the hierarchy of callers and callees. At the same time, we introduce static and dynamic annotation modules to supplement the lack of pseudocode semantics caused by stripping. In addition, in the above process, we iteratively add such annotations to maintain the semantic links between callers and callees. More details are provided in Section IV-A and IV-B.
@@
Conducting Evaluation: Previous work [24] indicated that existing metrics for generation tasks, such as Bilingual Evaluation Understudy (BLEU) [25], Metric for Evaluation of Translation with Explicit ORdering (METEOR) [26], Recall-Oriented Understudy for Gisting Evaluation-Longest Common Subsequence (ROUGE-L) [27], may not well-fit for evaluation of code summarization task. We thus employ BLEURT-sum, which is more sensitive to the quality of the pseudocode summary, thereby benefiting the evaluation in practice. More descriptions are given in Section V-A.
The contribution of this paper can be summarized as follows.
@@
  • A binary malware summarization Framework. We introduce Malsight, a novel framework that iteratively generates descriptions of binary malware, addressing challenges like entangled logic and stripped semantics.
@@
  • Large-scale datasets for binary malware summarization. We propose MalS and MalP, two novel datasets that can be used for the LLM training and testing of an LLM of binary malware summarization. To the best of our knowledge, the two datasets are the first in the field.
@@
  • An LLM-based binary malware summarization model. We propose MalT5, a lightweight (0.7B parameters) LLM tailored for binary malware summarization.
@@
  • An evaluation metric for the task: We present BLEURT-sum, a novel evaluation metric more sensitive to pseudocode summarization quality.
II Related Work
II-A Malware Analysis Engineering
@@
The field of malware analysis engineering focuses on analyzing malware’s functionality by examining its binaries, typically through static analysis methods that involve observing assembly code or pseudocode [28].
II-A1 Binary Decompilation
@@
Decompilation [29] converts executable files into human-readable pseudocode [30], which is more concise and structured than disassembled assembly code. Unlike disassembly, which maps instruction encoding directly to assembly statements, decompilation relies on algorithms and patterns (e.g. R2 [31], IDA [15], Ghidra [16]) and emerging methods using LLMs [32]. However, pseudocode lacks semantic information, such as function names. Decompiled function names are often unreadable (e.g., sub_4061C0 in IDA Pro) [33], providing limited pieces of information for further analysis.
II-A2 Human Static Analysis
@@
II-B NLP Technologies
In the Malsight process, we use Bidirectional Encoder Representation from Transformers (BERT) [37] and Text-to-Text Transfer Transformer (T5) [38] architecture language models to complete specific tasks.
II-B1 BERT Family
BERT, a transformer-based model pre-trained with Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), excels in Sequence Labeling (SL) tasks like Named Entity Recognition (NER). CodeBERT [39], extending BERT, learns code semantics through Code-Conditioned MLM and natural language generation, showing strong performance in code tasks. Given its BERT origins, we explore CodeBERT for Sequence Labeling (SL) in pseudocode.
II-B2 T5 Family
T5 [38] is a sequence-to-sequence Transformer model that unifies NLP tasks into a single framework. CodeT5 [40], built on T5, extends its capabilities to code understanding and generation using NL-PL bimodal data with identifier tagging and masked identifier prediction. CodeT5+ [41] improves on this with instruction tuning and additional pre-training tasks, enhancing performance on code-related tasks. Previous works like HexT5 [21] and BinT5 [20] further demonstrate the potential of T5-based models in binary code summarization for malware analysis.
II-B3 Transfer Learning
Transfer learning leverages abundant labelled data from a source task to help a model learn general features and supplement insufficient training datasets with similar or related data [42]. In Malsight, we fine-tune the CodeT5+ [41] model to achieve transfer learning from the source code summarization task to the decompiled code summarization task.
II-C Code Summary Evaluation
In code summary model evaluation, NLP text similarity algorithms, including word overlap and word embedding measures, compare generated results with reference sets, offering a cost-effective alternative to human evaluation.
II-C1 Words’ Overlap Measure
@@
II-C2 Words’ Embedding Measure
@@
Figure 2: Workflow of Malsight. The workflow takes binary malware as input and, after going through graph traversal, annotation generation, and model summarization, iteratively produces function-level code summaries based on the function call hierarchy.
III Motivation and Overview
The construction of the code summary framework mainly includes three steps, as shown in Fig. 2. During the evaluation phase, we test several evaluation methods and find a reasonable way to build an evaluation model for the code summary task. Simultaneously, our work involves the construction of multiple datasets (for subsequent stages of training and evaluation of MalT5).
III-A Code Summarization Process
@@
III-A1 Function List Extraction
Unlike existing code summarization methods focusing solely on a function’s internal information, we incorporate the call relationships between functions and treat the entire binary as a processing unit. As Fig. 2 shows, it is difficult for the subsequent code summarization model to correctly summarize the functionality of func_E without any information about func_F. (We assume that the function name of func_F has been corrupted.) Constructing a reverse-ordered function list provides a basis for the subsequent recovery of sub-functions functionality.
III-A2 Annotation Generation
Malsight will iteratively process the functions in the order of the reverse function list, adding annotations to each function. In other words, the program sequentially restores functions according to the Function Call Graph (FCG) diagram from the outermost to the innermost and passes function summary results inward.
III-A3 Code LLM Summary
Fun_E(annotated) is then fed into the code summary model for final code summary generation. In our work, we use transfer learning to adapt the model to both the malware functions’ functionality and the decompiled pseudocode’s structural features. The tokenizer splits the code into tokens and embeddings, incorporating a self-attention mechanism into a complete vector in the encoder. The decoder outputs a fine-tuned prediction based on the code summary and passes it back to the dynamic annotator for subsequent use.
@@ III-B Evaluation Method
We observe that the current popular text similarity measures make it difficult to evaluate the actual quality of the code summary objectively. Taking Fig. 3 as an example, two examples show the evaluation results of BLEU, Meteor, and ROUGE-L on two pairs of real code summaries. The figure shows that two code summaries without any semantically related results in high evaluation scores (blue-framed), while two semantically similar code abstracts receive low scores (red-framed), demonstrating the shortage of existing methods. In subsequent work, we propose BLEURT-sum, a BLEURT-based fine-tuning algorithm designed and evaluated to be more suitable for distinguishing available code summaries.
III-C Datasets Construction
@@
For the binary code summary and code summary model evaluation tasks, we build corresponding datasets.
III-C1 Dataset For Code Summary Model
To prevent data shift, training a code summary model requires a large dataset of malware pseudocode summaries. However, malware datasets [49] are usually available only as compiled binaries, often stripped and structurally complex, making it difficult to create summary datasets without labor-intensive manual efforts.